Since the variable exists on both sides, I need to use addition or subtraction to collect the variable on one side and get what 'x' equals.
6x - 2 = x + 13

Answer by checkley79(3341)

6x - 2 = x + 13
6x-x=13+2
5x=15
x=15/5
x=3 ans.
Proof:
6*3-2=3+13
18-2=16
16=16